AH+HA=DAD
each different letter in the two digit and three digit numbers represents a different digit. find the number represented by DAD

Answer by ankor@dixie-net.com(22740)

AH+HA=DAD
each different letter in the two digit and three digit numbers represents a different digit. find the number represented by DAD
:
Write it as to show the value of the numbers
(10A+H) + (10H+A) = 100D + 10A + D
Combine like terms
10A + A - 10A + H + 10H = 100D + D
A + 11H = 101D
A = 101D - 11H
Since we are dealing in single digit numbers, D has to equal 1, and H equal 9
A = 101(1) - 11(9)
A = 101 - 99
A = 2
Substituting, and checking
29 + 92 = DAD
121 = DAD